OpenStudy (anonymous):

Multiply (4.7 x 104) by (6.4 x 1015). Express your answer in scientific notation. A) 3.008 x 1020 B) 3.008 x 1060 C) 3.008 x 1061 Eliminate D) 30.08 x 1019

OpenStudy (mathstudent55):

\(\large (4.7 \times 10^4) \times (6.4 \times 10^{15}) \) First, use the commutative and associative properties to rearrange the multiplication this way: \(\large =(4.7 \times 6.4) \times (10^4 \times 10^{15} ) \) Ok so far?

OpenStudy (anonymous):

yah

OpenStudy (mathstudent55):

Good. Now multiply the first parentheses and multiply the second parentheses. What do you get?

OpenStudy (mathstudent55):

I have to go, so I'll explain the rest, and you can go over it and ask questions later.

OpenStudy (mathstudent55):

Now we multiply the numbers in each set of parentheses. \(\large = 30.08 \times 10^{4 + 15} \) \(\large =30.08 \times 10^{19} \) The multiplication was done correctly, but the answer above is not in scientific notation. In order to have scientific notation, the number has to be less than 10. We divide 30.08 by 10 and multiply the power of 10 by 10 \(\large = 3.008 \times 10^{20} \)

OpenStudy (anonymous):

okay thanks